How do you find a line perpendicular to 2x-5y=20?
Answer by stanbon(75887)

How do you find a line perpendicular to 2x-5y=20?
---
Find the slope of the given line:
5y = 2x-20
y = (2/5)x - 4
---
The the given line has slope = 2/5
------------------------------------------
The slope of any line perpendicular to the given line
must have a slope of -5/2
==================================
Cheers,
Stan H.